can i have an expression between particle velocity and wave velocity

Grade:11

1 Answers

Askiitians Expert Bharath-IITD
14 years ago

Dear Abhinav,

The concept of particle velocity and wave velocities come in case of a wave packet or group wave....

                 The particle velocity is nothing but the group velocity vg=dω/dk. And wave velocity is v=ω/k

which implies  ω = v*k substituting this we get  vg=( dv/dk )* k + v

 

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best  !!!

 


Regards,

Askiitians Experts

Adapa Bharath